# Quantum synchronization between two strongly driven YIG spheres mediated via a microwave cavity

### v1

@@ III Results
In this section, we will discuss the degree of quantum synchronization between two magnon modes by numerically solving the Eqs. (11), (12) and (II.1). We consider a strong driving regime, when the Rabi frequencies $\Omega_{i}$ and $\Omega_{c}$ of the driving fields are much larger than the decay rates $\gamma_{i}$ of the $i$th magnon mode and $\gamma_{c}$ of the the cavity mode, respectively. This regime would otherwise render an independent dynamics of the magnon mode in absence of the cavity mode, exhibiting an underdamped oscillation that could sustain upto a time scale $t\sim 1/\gamma_{i}$. In the present case, the cavity mode creates an indirect coupling between the magnon modes, leading to a classical synchronization between them that sustains up to a time-scale much larger than $1/\gamma_{i}$. We show these results in Figs. 2.
@@
We first choose $\Omega_{1}=1$ and $\Delta\Omega=\Omega_{2}-\Omega_{1}=10^{-5}$, for which both the magnon modes rapidly converge to a shared limit cycle in phase space (Figs. 2a2b, and 2c). This behavior is a clear signature of classical synchronization, where the mean quadratures $\bar{q}_{1}$, $\bar{q}_{2}$ and $\bar{p}_{1}$, $\bar{p}_{2}$ become indistinguishable at long times even when $t\approx 10^{5}/\Omega_{i}\gg 1/\gamma_{i}$. Such phase-locked dynamics of the vibrational degrees of freedom of two YIG spheres has been previously explored in a cavity-magnomechanical set up PhysRevResearch.5.043197 . When the Rabi frequency $\Omega_{2}$ is increased so that $\Omega_{2}-\Omega_{1}=0.1$, the system still maintains the phase synchronization, however with a nonzero phase difference $\phi=0.1320$ radian (Figs. 2d2e, and 2f). We note that such a phase relationship between coupled magnonic oscillators can be continuously tuned by varying coupling strengths PhysRevResearch.5.043197 ; PhysRevResearch.6.033207 .
@@
To quantify quantum synchronization, we next compute the $S_{q}^{\phi}$. Our results show that $S_{q}^{\phi}$ stabilizes at unity ($S_{q}^{\phi}=1$) at long times, indicating complete quantum synchronization between the magnon modes (see Fig. 3a). We have observed that the optimal value of $S_{q}^{\phi}$ remains unity for a wide range of the relative difference of the coupling constant $\Delta\Omega/\Omega_{1}$.
@@ III.1 Experimental consideration
Our parameter choices reflect realistic regimes established in cavity magnonics and magnomechanics, readily available to demonstrate our results. Normalized frequencies ($\Omega_{1,2}$,$\Omega_{c}\sim 1$) correspond to 1 GHz, typical for YIG-based systems zhang2016cavity ; li2019entangling . The small detuning ($\Omega_{2}-\Omega_{1}=10^{-5}$) is achievable via local magnetic field tuning li2019entangling . Coupling strengths ($g_{1,2}=0.1$) represent the strong-coupling regime (10–100 MHz), while damping rates ($\gamma_{1,2,c}=0.1$) match high-quality YIG spheres and microwave cavities zhang2016cavity . Kerr nonlinearities ($K_{1,2}=10^{-10}$) refer to weak intrinsic nonlinearity, consistent with low excitation levels wang2016magnon . These parameters enable robust synchronization and align with recent experimental and theoretical studies PhysRevResearch.5.043197 ; li2019entangling , grounded in well-established coupling mechanisms kittel1958interaction .
@@ IV Conclusion
In this work, we presented a comprehensive theoretical framework to investigate the synchronization between two magnon modes mediated by a single-mode microwave cavity. The magnon modes do not interact directly but are indirectly coupled through their mutual interaction with the cavity mode, which induces an effective nonlinear coupling. This nonlinear interaction, combined with intrinsic dissipation in the system, gives rise to complex dynamical behavior. Our analysis reveals that magnon modes can exhibit two distinct forms of synchronization. First, we observed classical synchronization (Fig. 2), where the dynamical trajectories of the magnon modes become phase-locked and evolve coherently and regularly over time.
@@ V ACKNOWLEDGMENTS
$\qquad-g_{i}A_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-C_{i})}-A_{i}\Omega_{i}\alpha_{i}^{*}e^{\iota t(D_{i}+C_{i})}\left(1+A_{i}|\alpha_{i}|^{2}\right)$
$\qquad+A_{i}\Omega_{i}\alpha_{i}e^{-\iota t(D_{i}+C_{i})}\left(2-A_{i}|\alpha_{i}|^{2}\right)$
$\qquad-A_{i}g_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-C_{i})}\left(1+A_{i}|\alpha_{i}|^{2}\right)\Bigg{]}\;$
@@
$\qquad-A_{i}\Omega_{i}\alpha_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(D_{i}+C_{i})}-A_{i}^{2}\Omega_{i}\alpha_{i}^{3}e^{-\iota t(D_{i}+C_{i})}$
$\qquad-A_{i}g_{i}\alpha_{i}\beta\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-C_{i})}\Bigg{]}\;$
$R_{i}$ $=-\iota g_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-C_{i})}\;$
$S_{i}$ $=\iota g_{i}A_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-C_{i})}\;$
$U_{i}$ $=-\iota g_{i}\left(1-A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(B_{i}-C_{i})}\;$
$W_{i}$ $=\iota A_{i}g_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-C_{i})}\;$
$T$ $=-\dfrac{\gamma_{c}}{2}\;$
$F_{i}$ $=\iota\Bigg{[}A_{i}g_{i}\alpha_{i}^{2}\beta^{*}e^{\iota t(B_{i}-C_{i})}-\Omega_{i}e^{\iota t(D_{i}+C_{i})}-g_{i}\beta e^{-\iota t(B_{i}-C_{i})}$
$\quad-A_{i}\Omega_{i}|\alpha_{i}|^{2}e^{\iota t(D_{i}+C_{i})}+A_{i}\Omega_{i}|\alpha_{i}|^{2}e^{-\iota t(D_{i}+C_{i})}$
$\quad-A_{i}g_{i}|\alpha_{i}|^{2}\beta e^{-\iota t(B_{i}-C_{i})}\Bigg{]}\;$
$F_{3}$ $=-\iota\Bigg{[}g_{1}e^{\iota t(B_{1}-C_{1})}-\iota\Omega_{c}e^{\iota t\Delta_{c}}\Bigg{]}\;$

### v2

@@ III Results
In this section, we will discuss by numerically solving the Eqs. (11), (12), and (II.1), how the quantum synchronization between two magnon modes is developed. We consider a strong driving regime, when the Rabi frequencies $\Omega_{i}$ and $\Omega_{c}$ of the driving fields are much larger than the decay rates $\gamma_{i}$ of the $i$th magnon mode and $\gamma_{c}$ of the cavity mode, respectively. This regime would otherwise render an independent dynamics of the magnon mode in the absence of the cavity mode, exhibiting an underdamped oscillation that could sustain up to a time scale $t\sim 1/\gamma_{i}$. In the present case, the cavity mode creates an indirect coupling between the magnon modes, leading to classical synchronization between them that persists up to a timescale much larger than $1/\gamma_{i}$. We show these results in Figs. 2.
@@
We first choose $\Delta\Omega=\Omega_{2}-\Omega_{1}=10^{-5}\Omega_{1}$, for which both the magnon modes rapidly converge to a shared limit cycle in phase space (Figs. 2a2b, and 2c). This behavior is a clear signature of classical synchronization, where the mean quadratures $\bar{q}_{1}$, $\bar{q}_{2}$ and $\bar{p}_{1}$, $\bar{p}_{2}$ become indistinguishable at times as long as $t\approx 10^{5}/\Omega_{i}\gg 1/\gamma_{i}$. Such phase-locked dynamics of two YIG spheres have been previously explored in a cavity-magnomechanical setup, in which the phase synchronization has been achieved between the vibrational degrees of freedom, instead of magnon modes PhysRevResearch.5.043197. When the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, the system still maintains the phase synchronization, however with a nonzero phase difference $\phi=0.1320$ radian (Figs. 2d2e, and 2f). Note that at the steady state, though the quadrature variables vary with time, the phase $\phi$ remains constant.
We further consider that the coupling strengths $g_{i}$ of these magnon modes with the cavity are different from each other. This could arise when the YIG spheres are trapped inside the cavity in an asymmetrical configuration with respect to the central antinodes of the cavity. When we choose $g_{1}>g_{2}$, the amplitude of the second magnon mode becomes lower compared to that of the first. Nevertheless, the system maintains phase synchronization with a phase difference of $\phi=0.1320$ radian (Fig. 2g, 2h and 2i). We note that such a phase relationship between coupled magnonic oscillators can be continuously tuned by varying coupling strengths PhysRevResearch.5.043197; PhysRevResearch.6.033207.
One of the important features of classical synchronization is its robustness against a large change in initial conditions. We verify the existence of limit cycles by varying the initial conditions (Fig. 3a, 3b, and 3c). We choose an initial condition $\bar{q}_{1}(0),\bar{p}_{1}(0)]=[1,0]$ and $\bar{q}_{2}(0),\bar{p}_{2}(0)]=[2,0]$. For $\Delta\Omega=10^{-5}\Omega_{1}$, both magnon modes rapidly converge to a common limit cycle in phase space. Similar feature persists also when the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, though with a finite phase difference of $\phi=1.2276$ radian (Figs. 3d, 3e, and 3f). In this regime, the system exhibits chaotic transients before settling into a shared limit cycle, providing a clear signature of synchronization. Furthermore, when the detunings of the driving fields are chosen differently, e.g., $\Delta_{1}=0.001$ and $\Delta_{2}=0.0025$, the system continues to display phase synchronization with an increased phase difference of $\phi=1.5337$ radian (Figs. 3g, 3h, and 3i).
@@
To quantify quantum synchronization, we next compute the $S_{q}^{\phi}$. Our results show that $S_{q}^{\phi}$ stabilizes at unity ($S_{q}^{\phi}=1$) at long times, indicating complete quantum synchronization between the magnon modes (see Fig. 4a). We have observed that the optimal value of $S_{q}^{\phi}$ remains unity for a wide range of the relative difference of the coupling constant $\Delta\Omega/\Omega_{1}$ and of the detunings $(\Delta_{2}-\Delta_{1})$. We have verified that $S_{q}^{\phi}$ remains unity at long times, even when the two magnon modes start from different initial conditions (results not shown). It is interesting to note that the $S_{q}^{\phi}$ becomes unity at a time scale $t\sim 50/\Omega_{1}$, which is much larger than the decay time-scale $1/\gamma_{1}\sim 10/\Omega_{1}$.
@@ III.1 Experimental consideration
Our parameter choices reflect realistic regimes established in cavity magnonics and magnomechanics, readily available to demonstrate our results. Normalized frequencies ($\Omega_{1,2}$$\Omega_{c}\sim 1$) correspond to 1 GHz, typical for YIG-based systems zhang2016cavity; li2019entangling. The small detuning ($\Delta\Omega=10^{-5}\Omega_{1}$) is achievable via local magnetic field tuning li2019entangling. Coupling strengths ($g_{1,2}=0.1$) represent the strong-coupling regime $g_{1}\sim\gamma_{i},\gamma_{c}$ (10–100 MHz), while damping rates ($\gamma_{1,2,c}=0.1$) match high-quality YIG spheres and microwave cavities zhang2016cavity. Kerr nonlinearities ($K_{1,2}=10^{-10}$) refer to weak intrinsic nonlinearity, consistent with low excitation levels wang2016magnon. These parameters enable robust synchronization and align with recent experimental and theoretical studies PhysRevResearch.5.043197; li2019entangling, grounded in well-established coupling mechanisms kittel1958interaction.
@@ IV Conclusion
In this work, we presented a comprehensive theoretical framework to investigate the synchronization between two magnon modes mediated by a single-mode microwave cavity. The magnon modes do not interact directly but are indirectly coupled through their mutual interaction with the cavity mode, which induces an effective nonlinear coupling. This nonlinear interaction, combined with intrinsic dissipation in the system, gives rise to complex dynamical behavior. Our analysis reveals that magnon modes can exhibit two distinct forms of synchronization. First, we observed classical synchronization with different initial conditions (Fig. 2) and (Fig. 3), where the dynamical trajectories of the magnon modes become phase-locked and evolve coherently and regularly over time.
@@ V ACKNOWLEDGMENTS
$\hskip 17.00024pt-g_{i}A_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-C_{i})}-A_{i}\Omega_{i}\alpha_{i}^{*}e^{\iota t(D_{i}+C_{i})}\left(1+A_{i}|\alpha_{i}|^{2}\right)$
$\hskip 17.00024pt+A_{i}\Omega_{i}\alpha_{i}e^{-\iota t(D_{i}+C_{i})}\left(2-A_{i}|\alpha_{i}|^{2}\right)$
$\hskip 17.00024pt-A_{i}g_{i}\alpha_{i}^{*}\beta e^{-\iota t(B_{i}-C_{i})}\left(1+A_{i}|\alpha_{i}|^{2}\right)\Bigg]\;,$
@@
$\hskip 17.00024pt-A_{i}\Omega_{i}\alpha_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(D_{i}+C_{i})}-A_{i}^{2}\Omega_{i}\alpha_{i}^{3}e^{-\iota t(D_{i}+C_{i})}$
$\hskip 17.00024pt-A_{i}g_{i}\alpha_{i}\beta\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-C_{i})}\Bigg]\;,$
$R_{i}$ $=-\iota g_{i}\left(1+A_{i}|\alpha_{i}|^{2}\right)e^{-\iota t(B_{i}-C_{i})}\;,$
$S_{i}$ $=\iota g_{i}A_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-C_{i})}\;,$
$U_{i}$ $=-\iota g_{i}\left(1-A_{i}|\alpha_{i}|^{2}\right)e^{\iota t(B_{i}-C_{i})}\;,$
$W_{i}$ $=\iota A_{i}g_{i}\alpha_{i}^{2}e^{\iota t(B_{i}-C_{i})}\;,$
$T$ $=-\dfrac{\gamma_{c}}{2}\;,$